Calculus I, Fall 2014 (http://www.math.nthu.edu.tw/~wangwc/)

Brief answer to selected problems in Homework 16

1. Section 8.4 Problem 32: ∫ \frac{θ^{4}−4θ^{3}+2θ^{2}−3θ+1}{(θ^{2}+1)^{3}}dθ = ∫ \frac{1}{θ^{2}+1} − \frac{4θ}{(θ^{2}+1)^{2}} + \frac{θ}{(θ^{2}+1)^{3}} dθ